OpenStudy (anonymous):

ahhh, negative exponent: 2^-4 + 2^-6 / (2^-3) = ?

OpenStudy (anonymous):

2^-4 + 2^-6 / (2^-3) = 1/2^4 + 1/2^6 / 1/2^3 = 2^2+1 / 2^6 / 1/2^3

OpenStudy (anonymous):

= 2^2 + 1/2^6 x 2^3 / 1 = 2^2+1/2^3 = 5/8

OpenStudy (anonymous):

Why is the second step: 2^2+1 / 2^6 / (1/2^3)? omg I'm still so confused

OpenStudy (anonymous):

i thought I got it but ahh.

sam (.sam.):

\[\frac{2^{-4} + 2^{-6} }{ 2^{-3}}?\]

OpenStudy (callisto):

If the answer is 5/8, then it's \[\frac{2^{-4} + 2^{-6} }{ 2^{-3}}\] \[\frac{2^{-4} + 2^{-6} }{ 2^{-3}}\]\[\large = \frac{\frac{1}{2^{4}} + \frac{1}{2^{6}} }{ \frac{1}{2^{3}}}\]\[\large = \frac{\frac{2^2}{2^{4+2}} + \frac{1}{2^{6}} }{ \frac{1}{2^{3}}}\]\[\large = \frac{\frac{2^2+1}{2^{6}} }{ \frac{1}{2^{3}}}\]\[\large = \frac{\frac{5}{2^{6}} }{ \frac{1}{2^{3}}}\]\[ = 2^3(\frac{5}{2^{6}}) \]\[= \frac{5}{2^{6-3}} \]\[= \frac{5}{2^{3}} \]\[= \frac{5}{8} \]

OpenStudy (anonymous):

why is it 2^2 for second step? :$

OpenStudy (callisto):

Multiply both denominator and numerator by 2^2 so that you can get 2^6 for the denominator and do the addition for the fractions

OpenStudy (anonymous):

wait so first it's 1/2^4 + 1/2^6 and so why is it now 2^2? I subtract exponent 6 from exponent 4?

OpenStudy (callisto):

\[\frac{1}{2^4} + \frac{1}{2^6} = \frac{1\times 2^2}{2^4 \times 2^2} + \frac{1}{2^6} = \frac{1\times 2^2}{2^{4+2}} + \frac{1}{2^6} = \frac{2^2}{2^{6}} + \frac{1}{2^6}\]Got it?

OpenStudy (anonymous):

i'm so sorry but why 1 x 2^2?

OpenStudy (callisto):

Because you've multiply the denominator by 2^2, you can't change the value of the fraction. So, you need to do the same for numerator

OpenStudy (anonymous):

ohh so for the denominator you multiplied 2^4 x 2^2 to get 2^6?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

thank you!

OpenStudy (callisto):

Are you sure you understand this time?

OpenStudy (anonymous):

yes, now I'm sure. Thank you so much!!

OpenStudy (callisto):

Welcome :) Special thanks to @.Sam.

OpenStudy (anonymous):

yes, thank you to @.Sam.